GRAVITATON WAVE AND GRAVITATIONAL-PHOTON INTERACTION

\underline{KHOLMURAD KHASANOV}, Lomonosov Moscow State University, Gas and Wave Dynamics Department, Moscow, Russia, 119991, GSP-1, 1 Leninskie Gory Str.Email: kholkh@bk.ru.

Gravitation waves and gravitational-photon interaction with high energy photons emission is found experimentally. Supercompressibility phenomenon was studied. Spectral investigations of supersonic jets and incandescent nichrome thread and wolfram spiral were studied. The shifting of the emission spectrum was detected depending on vector of gravity. The increasing frequency of light emitted against gravity vector is measured. Uneven along the spectrum character of intensity increasing is found. Generation of short-wavelength component of the spectrum is observed in case of more power of heating. The measurements show that presented interactions have resonance nature. Our experiments demonstrate the existence resonance nature. Our experiments demonstrate the gravitation wave and generation and existence of gravitational-photon interactions.

From left to right: Fig. 1-2. Visualization of the gravitation wave. Fig. 3-5. Gravitational-photon interaction in HF field.